# Enhancing Real-World Complex Network Representations with Hyperedge Augmentation

### v1

@@ Abstract.
Graph augmentation methods play a crucial role in improving the performance and enhancing generalisation capabilities in Graph Neural Networks (GNNs). Existing graph augmentation methods mainly perturb the graph structures and are usually limited to pairwise node relations. These methods cannot fully address the complexities of real-world large-scale networks that often involve higher-order node relations beyond only being pairwise. Meanwhile, real-world graph datasets are predominantly modelled as simple graphs, due to the scarcity of data that can be used to form higher-order edges. Therefore, reconfiguring the higher-order edges as an integration into graph augmentation strategies lights up a promising research path to address the aforementioned issues. In this paper, we present Hyperedge Augmentation (HyperAug), a novel graph augmentation method that constructs virtual hyperedges directly form the raw data, and produces auxiliary node features by extracting from the virtual hyperedge information, which are used for enhancing GNN performances on downstream tasks. We design three diverse virtual hyperedge construction strategies to accompany the augmentation scheme: (1) via graph statistics, (2) from multiple data perspectives, and (3) utilising multi-modality. Furthermore, to facilitate HyperAug evaluation, we provide 23 novel real-world graph datasets across various domains including social media, biology, and e-commerce. Our empirical study shows that HyperAug consistently and significantly outperforms GNN baselines and other graph augmentation methods, across a variety of application contexts, which clearly indicates that it can effectively incorporate higher-order node relations into graph augmentation methods for real-world complex networks. Specifically, on the Amazon datasets, our method surpass the GCN baseline by an average of 37.0%, outperform the GAT baseline by an average of 30.8%, and exceed the GraphSAGE baseline by an average of 33.8%.
Graph Neural Networks, Complex Networks, Data Augmentation, Multimodal Learning
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: arXiv preprint; August 25–29, 2024; Barcelona, Spainisbn: 978-X-XXXX-XXXX-X/YY/MMccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Neural networksccs: Information systems Information retrievalccs: Information systems Data mining
1. Introduction
@@
Figure 1. An overview of the HyperAug method showcasing three virtual hyperedge construction strategies: via graph statistics, from multiple data perspectives, and using multiple data modalities. The process initiates with an original graph $\mathcal{G}$. The virtual hyperedge construction methods contribute to the construction of an augmented graph $\mathcal{G^{\prime}}$, which is then processed through a feature extraction function $f$ to obtain a set of auxiliary node features $\mathbf{Z}^{\prime}$. The auxiliary features $\mathbf{Z}^{\prime}$ are then combined with the node embeddings $\mathbf{Z}^{\prime}$ obtained from the original graph, regularised by a linear layer with weights $\bm{\theta}$ and bias $\mathbf{b}$, to produce the final augmented node embeddings, which are then used in downstream tasks to enhance prediction accuracy.
@@
Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Brody et al., 2022; Zeng et al., 2020) are powerful tools for learning the representation of relationships between objects, ranging from social networks (Monti et al., 2019), biology (Rozemberczki et al., 2021) to e-commerce (McAuley et al., 2015; He and McAuley, 2016; Ni et al., 2019). Representation learning tasks, such as node prediction, constitute a major category of tasks for Graph Neural Networks (GNNs). However, limitations in the generalisability of GNNs have obstructed their broader application in real-world settings (Rusch et al., 2023).
@@
Several graph augmentation techniques have been introduced to enhance the performance and foster better generalisation for GNNs, specifically in the context of node prediction tasks invovling large graphs. Current graph augmentation methods include graph structures perturbation (Rong et al., 2020; Ding et al., 2022), feature perturbation (Feng et al., 2019a), and label-oriented augmentations (Han et al., 2022; Wang et al., 2021). Simple-GNNs – GNNs that operate on simple graphs with only pairwise relations, normally use perturbation-based augmentation techniques. However, these augmentation methods do not adequately grasp the complexities of real-world, large-scale networks that frequently display intricate node relationships beyond having only pairwise connections. While Hypergraph Neural Networks (hyper-GNNs) (Yadati et al., 2019; Dong et al., 2020; Wang et al., 2023) have been developed to model higher-order node relations, they encounter a significant challenge due to the scarcity of explicitly recorded data to form high-order edges in many real-world datasets. This scarcity hampers the direct application of hyper-GNNs, as they rely on comprehensive data that captures complex node relations.
@@
Although the scarcity of higher-order edges presents a challenge, integrating advanced modeling methods into augmentation strategies offers a promising avenue to address this issue. Traditionally, graph augmentation has focused on first-order connections, primarily modifying or adding direct links between nodes (Rong et al., 2020; Ding et al., 2022). However, this approach overlooks the rich, multi-level interactions captured in higher-order edges, which encompass indirect connections, such as patterns of group interactions within the networks. By incorporating these higher-order relations, graph augmentation can achieve a more nuanced understanding of network dynamics, revealing hidden structures and patterns that are not apparent in direct links alone. This approach is particularly beneficial in real-world complex networks, where the interplay of various types of connections can provide deeper insights into the underlying system, across a wide range of domains including social media, biological ecosystems and e-commerce network. The integration of higher-order edge information into graph augmentation strategies therefore opens up new possibilities for more sophisticated and accurate network analysis.
@@
In our study, we introduce a suite of graph augmentation techniques that extend beyond conventional approaches. Our methodology encompasses higher-order node relations, and we present a novel graph augmentation strategy termed Hyperedge Augmentation (HyperAug). As illustrated in Figure 1, HyperAug employs a two-step augmentation mechanism. First, HyperAug constructs virtual hyperedges from the original graph. Second, HyperAug further processes the virtual hyperedges through a feature extraction function ($f$ in Figure 1) to produce auxiliary node embeddings. These auxiliary embeddings can then be used by GNNs to improve their robustness and accuracy on downstream tasks. We present three distinct ways for constructing these virtual hyperedges: (1) via graph statistics, (2) from multiple data perspectives, and (3) using multiple data modalities. Our augmentation schemes showcase notable flexibility across a variety of datasets, catering to diverse applications such as social, biological, e-commerce, and knowledge networks. Our main contributions are as follows:
@@
  • We introduce HyperAug, a novel data augmentation method for GNNs on node prediction tasks. HyperAug allows the network to integrate high-order edge information. Along with our method, we also release 23 new datasets, encompassing diverse domains including social media, biological networks as well as e-commerce graphs.
@@
  • We assess various design choices within HyperAug, including different methods for constructing virtual hyperedges and varying feature extraction functions. Our examination of diverse HyperAug variants across multiple GNNs and prediction tasks underscores HyperAug’s versatility for various application contexts.
@@
  • Through a comprehensive empirical evaluation on various real-world datasets and GNN architectures, we demonstrate that HyperAug offers consistent and significant performance improvements. For example, on the Amazon datasets, HyperAug boosts the node prediction accuracy of GCN (Kipf and Welling, 2017) by an average of 37.0%, GAT (Veličković et al., 2018) by an average of 30.8%, and GraphSAGE (Hamilton et al., 2017) by an average of 33.8%.
2. Related Work
@@
Graph Neural Networks (GNNs) on simple graphs encode the nodes through neural networks and learn the representations of the nodes through message-passing within the graph structure. GCN incorporates the convolution operation in computer vision into GNNs (Kipf and Welling, 2017) . GAT (Veličković et al., 2018) and GATv2 (Brody et al., 2022) are another family of GNN variants that focus on improving the expressive power of GNNs by using the attention mechanism. GraphSAGE (Hamilton et al., 2017) is a general inductive framework that leverages node information to efficiently generate node embeddings for previously unseen data. GraphSAINT (Zeng et al., 2020) underscores the significance of graph sampling-based inductive learning – it utilises diverse graph sampling techniques and illustrates how learning on smaller, sampled graphs can enhance training efficiency, particularly with large graphs. While these models demonstrate the success on simple graph datasets, they continue to face challenges in their generalisation to unseen data and are susceptible to small variations in graph structures (Cai and Wang, 2020).
@@
Enhancing the generalisability of GNNs through graph augmentation methods remains a key area of research interest in recent years. Existing graph augmentation methods introduce perturbations to the graph structure, enabling GNNs trained on these altered graphs to learn and capture invariances. This includes adding perturbations to the graphs’ adjacency matrices through DropEdge (Rong et al., 2020), randomly removing nodes through DropNode (You et al., 2020; Ding et al., 2022), and perturbing node edge features through feature masking. Another line of graph augmentations relies on the generation of synthetic data, which can be acheived through the interpolation of existing data via Mixup (Zhang et al., 2018; Han et al., 2022; Wang et al., 2021), or using generative models to enrich the training data (Liu et al., 2023). However, these existing augmentation strategies do not inherently increase the expressiveness of GNNs. Standard GNNs typically aggregate information from their neighbourhoods, a process that shares limitations with the Weisfeiler-Lehman (1-WL) graph isomorphism test (Weisfeiler and Leman, 1968; Xu et al., 2019; Hamilton, 2020). For instance, as depicted in Figure 2(a), simple-GNNs fail to distinguish between two structurally different graphs with the same node degrees due to this limitation. This research aims to develop augmentation-based methods that address and overcome these constraints, thereby enhancing the capabilities of GNNs beyond their current limits.
@@
Hypergraphs are designed to capture more complex node relations, where an edge can connect two or more nodes. In general, GNNs for hypergraphs optimise the node representation through a two-step process. Initially, the node embeddings within each hyperedge are aggregated to form a hidden embedding of each hyperedge. Subsequently, the hidden embeddings of hyperedges with common nodes are aggregated to update the representations of their common nodes. Both HGNN (Feng et al., 2019b) and HyperConv (Bai et al., 2021) precisely follow this process. The expressiveness of hypergraph GNNs could be enhanced by modifying this procedure. For instance, HyperGCN (Yadati et al., 2019) refines the node aggregation within hyperedges using mediators (Chan and Liang, 2020). HyperAtten (Bai et al., 2021) uses attention to measure the degree to which a node belongs to a hyperedge. HNHN (Dong et al., 2020) applies nonlinear functions to both node and edge aggregation processes. ED-HNN (Wang et al., 2023) approximates continuous equivariant hypergraph diffusion operators on hypergraphs by feeding node representations into the message from hyperedges to nodes. GNNs designed for hypergraphs are effective in modelling complex networks. However, real-world datasets are predominantly recorded as simple graphs, which has limited the application of hyper-GNNs due to the scarcity of data that can form hyperedges. In this study, we propose HyperAug that derives hyperedges directly from the raw data, thereby extending the applicability of hypergraph GNNs to cases where only conventional simple graph data is available. By utilising higher-order GNNs on virtually established hyperedges to produce auxiliary features, HyperAug effectively broadens the potential of hyperedges to scenarios previously confined to simple graph data.
3. Method
3.1. Hyperedge Construction
Table 1. Aggregated statistics of the datasets used for HyperAug evaluation. Full statistics can be found in Table 7 in Appendix A.
Name #Graphs Avg. #Nodes Avg. #Edges Avg. #Hyperedges Avg. Node Degree Avg. Hyperedge Degree #Classes
MUSAE-GitHub 1 37,700 578,006 223,672 30.7 4.6 4
MUSAE-Facebook 1 22,470 342,004 236,663 30.4 9.9 4
MUSAE-Twitch 6 5,686 143,038 110,142 50.6 6.0 2
MUSAE-Wiki 3 6,370 266,998 118,920 88.8 14.4 Regression
GRAND-Tissues 6 5,931 5,926 11,472 2.0 1.3 3
GRAND-Diseases 4 4,596 6,252 7,743 2.7 1.3 3
Cora 2 2,708 5,429 1,326 4.0 3.5 7
Pubmed 1 19,717 44,338 7,963 4.5 4.3 3
Amazon-Computers 1 10,226 55,324 10,226 10.8 4.0 10
Amazon-Photos 1 6,777 45,306 6,777 13.4 4.8 10
Formally, a simple graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a collection of nodes $\mathcal{V}$ and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ between pairs of nodes. This graph abstraction assumes that each edge only connects two nodes. However, as discussed in the previous sections, many real-world networks have more complex node relations than just pairwise relations. Inspired by the hypergraph modelling, HyperAug attempts to capture such complex relations, by incorporating virtual hyperedges—edges that can connect more than two nodes—into the network. Specifically, HyperAug augments a simple graph by constructing a set of virtual hyperedges $\mathcal{E}_{h}$ from the original graph, producing an augmented graph $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E},\mathcal{E}_{h})$. Each hyperedge $e_{h}\in\mathcal{E}_{h}$ is a non-empty subset of $\mathcal{V}$, which contains at least three nodes, thereby capturing complex multi-node relations. Depending on the nature of the original graph, the virtual hyperedges can be constructed in one of the following three ways: from the graph statistics, from a different data perspective, or from a different data modality.
From the Graph Statistics.
For graphs that can reveal valuable information from their local clusters and cliques, such as the social networks, HyperAug constructs virtual hyperedges by computing from graph statistics, such as maximal cliques. HyperAug then groups those nodes within the same maximal cliques. This virtual hyperedge augmentation acts as a complement to the simple edges in the original graph by calculating the local clustering relationships amongst the nodes. This approach enables GNNs seeking to utilise local clustering information to rapidly access that information without needing to explicitly pre-compute it. In addition, by separating out the local clusters and cliques from the simple edge adjacency of the original graph, HyperAug can prevent GNNs from mixing up those information, thereby potentially enhancing its training performance. Another crucial advantage for this type of augmentation is that it enables GNNs to overcome the expressiveness limitations posed by the Weisfeiler-Lehman (1-WL) graph isomorphism test (Weisfeiler and Leman, 1968; Xu et al., 2019; Hamilton, 2020): it is able to help GNNs distinguish non-isomorphic graphs with identical node degrees, by specifying their distinct maximal cliques, as illustrated in Figure 2.
(a) Without HyperAug
(b) With HyperAug

### v2

# Enhancing Node Representations for Real-World Complex Networks with Topological Augmentation

@@ Abstract
Graph augmentation methods play a crucial role in improving the performance and enhancing generalisation capabilities in Graph Neural Networks (GNNs). Existing graph augmentation methods mainly perturb the graph structures, and are usually limited to pairwise node relations. These methods cannot fully address the complexities of real-world large-scale networks, which often involve higher-order node relations beyond only being pairwise. Meanwhile, real-world graph datasets are predominantly modelled as simple graphs, due to the scarcity of data that can be used to form higher-order edges. Therefore, reconfiguring the higher-order edges as an integration into graph augmentation strategies lights up a promising research path to address the aforementioned issues. In this paper, we present Topological Augmentation (TopoAug), a novel graph augmentation method that builds a combinatorial complex from the original graph by constructing virtual hyperedges directly from the raw data. TopoAug then produces auxiliary node features by extracting information from the combinatorial complex, which are used for enhancing GNN performances on downstream tasks. We design three diverse virtual hyperedge construction strategies to accompany the construction of combinatorial complexes: (1) via graph statistics, (2) from multiple data perspectives, and (3) utilising multi-modality. Furthermore, to facilitate TopoAug evaluation, we provide 23 novel real-world graph datasets across various domains including social media, biology, and e-commerce. Our empirical study shows that TopoAug consistently and significantly outperforms GNN baselines and other graph augmentation methods, across a variety of application contexts, which clearly indicates that it can effectively incorporate higher-order node relations into the graph augmentation for real-world complex networks.
[…]
234
1 Introduction
@@
Figure 1: An overview of the TopoAug method showcasing three virtual hyperedge construction strategies: via graph statistics, from multiple data perspectives, and using multiple data modalities. The process initiates with an original graph $\mathcal{G}$. The virtual hyperedge construction methods contribute to the construction of an combinatorial complex $\mathcal{X}$, which is then processed through an auxiliary function $f_{\text{aux}}$ to obtain a set of auxiliary node features $\mathbf{Z}^{\prime}$. The auxiliary features $\mathbf{Z}^{\prime}$ are then combined with the unaugmented node embeddings $\mathbf{Z}$ (obtained from the original graph via the backbone embedding function $f_{\text{emb}}$), to produce the final augmented node embeddings $\mathbf{Z}_{\text{aug}}$. The final augmented node embeddings $\mathbf{Z}_{\text{aug}}$ are then used in downstream tasks to enhance prediction accuracy.
@@
Graph Neural Networks (GNNs) [3, 13, 20, 33, 41] are powerful tools for learning the representation of relationships between objects, ranging from social networks [25], biology [29] to e-commerce [16, 24, 26]. Representation learning tasks, such as node prediction, constitute a major category of tasks for GNNs. However, limitations in the generalisability of GNNs have obstructed their broader application in real-world settings [30].
@@
Several graph augmentation techniques have been introduced to enhance the performance and foster better generalisation for GNNs, specifically in the context of node prediction tasks involving large graphs. Current graph augmentation methods include graph structures perturbation [6, 28], feature perturbation [9], and label-oriented augmentations [15, 35]. Simple-GNNs – GNNs that operate on simple graphs with only pairwise relations – normally use perturbation-based augmentation techniques. However, these augmentation methods do not adequately grasp the complexities of real-world, large-scale networks that frequently display intricate node relationships beyond having only pairwise connections. While higher-order GNNs, such as Hypergraph Neural Networks (hyper-GNNs) [7, 34, 39], have been developed to model higher-order node relations, they encounter a significant challenge due to the scarcity of explicitly recorded data for forming high-order edges in many real-world datasets. This scarcity hampers the direct application of hyper-GNNs, as they rely on comprehensive data that capture complex node relations.
@@
Although the scarcity of higher-order edges presents a challenge, integrating advanced modeling methods into augmentation strategies offers a promising avenue to address this issue. Traditionally, graph augmentation has focused on first-order connections, primarily modifying or adding direct links between nodes [6, 28]. However, this approach overlooks the rich, multi-level interactions captured in higher-order edges, which encompass indirect connections, such as patterns of group interactions within the networks. By incorporating these higher-order relations, graph augmentation can achieve a more nuanced understanding of network dynamics, revealing hidden structures and patterns that are not apparent in direct links alone. This approach is particularly beneficial in real-world complex networks, where the interplay of various types of connections can provide deeper insights into the underlying system, across a wide range of domains including social media, biological ecosystems and e-commerce networks. Therefore, the integration of higher-order edge information into graph augmentation strategies opens up new possibilities for more sophisticated and accurate network analysis.
@@
In our study, we introduce a suite of graph augmentation techniques that extend beyond conventional approaches. Our methodology encompasses higher-order node relations, and we present a novel graph augmentation strategy termed Topological Augmentation (TopoAug). As illustrated in Figure 1, TopoAug employs a two-step augmentation mechanism. First, TopoAug constructs hyperedges from the original graph data, thereby forming a combinatorial complex [12]. Second, TopoAug further processes the combinatorial complex through an auxiliary function ($f_{\text{aux}}$ in Figure 1) to produce auxiliary node embeddings. These auxiliary embeddings can then be used by GNNs to improve their robustness and accuracy on downstream tasks. We present three distinct ways for constructing these hyperedges: (1) via graph statistics, (2) from multiple data perspectives, and (3) using multiple data modalities. Our augmentation schemes showcase notable flexibility across a variety of datasets, catering to diverse applications such as social, biological, e-commerce, and knowledge networks. Our main contributions are as follows:
@@
  • We introduce TopoAug, a novel data augmentation method for GNNs on node prediction tasks. TopoAug allows the network to integrate high-order edge information. Along with our method, we also release 23 new datasets, encompassing diverse domains including social media, biology as well as e-commerce networks.
@@
  • We assess various design choices within TopoAug, including different methods for constructing virtual hyperedges and varying auxiliary functions. Our examination of diverse TopoAug variants across multiple GNNs and prediction tasks underscores TopoAug’s versatility for various application contexts.
@@
  • Through a comprehensive empirical evaluation on various real-world datasets and GNN architectures, we demonstrate that Topo-Aug offers consistent and significant performance improvements over GNN baselines and existing graph augmentation methods.
2 Related Work
Topological Deep Learning
Topological deep learning extends machine learning models from graph data to data on topological domains, such as hypergraphs and other higher-order graphs, to facilitate the exponential growth in both the amount and the complexity of data for computational analysis. Formally, a (simple) graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a collection of nodes $\mathcal{V}$ and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ between pairs of nodes. This simple graph abstraction assumes that each edge only connects two nodes. However, as discussed in the previous sections, many real-world networks have more complex node relations than just pairwise relations. A hypergraph is also defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, but $\mathcal{E}\subseteq\mathscr{P}(\mathcal{V})\setminus\varnothing$ is now the set of hyperedges, each of which can connect two or more nodes. However, as real-world datasets are predominantly recorded as simple graphs, treating both simple edges and “true” hyperedges (edges that connects strictly more than two nodes) as hyperedges altogether also limits their flexibility, due to the scarcity of data that can form true hyperedges. Recently, there emerges a novel type of topological domain – combinatorial complexes, that enrich the hypergraphs by separating simple edges and true hyperedges through the inclusion of hierarchical ranks [12]:
Definition 1 (Combinatorial complex).
A combinatorial complex is a triple $(\mathcal{V},\mathcal{X},\mathrm{rk})$ consisting of a node set $\mathcal{V}$, a node relation set $\mathcal{X}\subseteq\mathscr{P}(\mathcal{V})\setminus\varnothing$, and a rank function $\mathrm{rk}:\mathcal{X}\to\mathbb{N}$, such that
  1. 1.
    $\forall v\in\mathcal{V}.\>\{v\}\in\mathcal{X}$; and
  2. 2.
    the rank function $\mathrm{rk}$ is order-preserving, which means that
    $\forall x,y\in\mathcal{X}.\>x\subseteq y\implies\mathrm{rk}(x)\leq\mathrm{rk}(y)$.
For brevity, $\mathcal{X}$ is used as an abbreviated notation for a combinatorial complex $(\mathcal{V},\mathcal{X},\mathrm{rk})$. Typically, the rank of any singleton node relation $\{v\}$ in $\mathcal{X}$ is set to zero, to make it naturally aligned with simple graphs and hypergraphs. The rank function effectively induces a hierarchical structure on $\mathcal{X}$, which can be used to separate simple edges and true hyperedges, by assigning them with different ranks.
@@ 2 Related Work
Graph Neural Networks (GNNs) on simple graphs encode the nodes through neural networks, and learn the representations of the nodes through message-passing within the graph structure. GCN incorporates the convolution operation in computer vision into GNNs [20]. GAT [33] and GATv2 [3] are another family of GNN variants that focus on improving the expressive power of GNNs by using the attention mechanism. GraphSAGE [13] is a general inductive framework that leverages node information to efficiently generate node embeddings for previously unseen data. GraphSAINT [41] underscores the significance of graph sampling-based inductive learning – it utilises diverse graph sampling techniques and illustrates how learning on smaller, sampled graphs can enhance training efficiency, particularly with large graphs. While these models demonstrate the success on simple graph datasets, they continue to face challenges in their generalisation to unseen data and are susceptible to small variations in graph structures [4].
@@
Enhancing the generalisability of GNNs through graph augmentation methods remains a key area of research interest in recent years. Existing graph augmentation methods introduce perturbations to the graph structure, enabling GNNs trained on these altered graphs to learn and capture invariance. This includes adding perturbations to the graphs’ adjacency matrices through DropEdge [28], randomly removing nodes through DropNode [6, 40], and perturbing node edge features through feature masking. Another line of graph augmentations relies on the generation of synthetic data, which can be achieved through the interpolation of existing data via Mixup [15, 35, 42], or using generative models to enrich the training data [21]. However, these existing augmentation strategies do not inherently increase the expressiveness of GNNs. Standard GNNs typically aggregate information from their neighbourhoods, a process that shares limitations with the Weisfeiler-Lehman (1-WL) graph isomorphism test [14, 37, 38]. For instance, as depicted in Figure 2(a), simple-GNNs fail to distinguish between two structurally different graphs with the same node degrees due to this limitation. This research aims to develop augmentation-based methods that address and overcome these constraints, thereby enhancing the capabilities of GNNs beyond their current limits.
@@
Hypergraphs are designed to capture more complex node relations, where an edge can connect two or more nodes. In general, GNNs for hypergraphs optimise the node representation through a two-step process. Initially, the node embeddings within each hyperedge are aggregated to form a hidden embedding of each hyperedge. Subsequently, the hidden embeddings of hyperedges with common nodes are aggregated to update the representations of their common nodes. Both HGNN [10] and HyperConv [1] precisely follow this process. The expressiveness of hypergraph GNNs could be enhanced by modifying this procedure. For instance, HyperGCN [39] refines the node aggregation within hyperedges using mediators [5]. HyperAtten [1] uses attention to measure the degree to which a node belongs to a hyperedge. HNHN [7] applies nonlinear functions to both node and edge aggregation processes. ED-HNN [34] approximates continuous equivariant hypergraph diffusion operators on hypergraphs by feeding node representations into the message from hyperedges to nodes. GNNs designed for hypergraphs are effective in modelling complex networks. However, real-world datasets are predominantly recorded as simple graphs, which has limited the application of hyper-GNNs due to the scarcity of data that can form hyperedges. In this study, we propose TopoAug that derives hyperedges directly from the raw data, thereby extending the applicability of hyper-GNNs to cases where only conventional simple graph data are available. By utilising higher-order GNNs on virtually established hyperedges to produce auxiliary features, TopoAug effectively broadens the potential of hyperedges to scenarios previously confined to simple graph data.
3 Method
3.1 Combinatorial Complex Construction
Inspired by the combinatorial complex modelling, TopoAug attempts to capture complex relations in real-world large networks by constructing a combinatorial complex from the original network. Specifically, TopoAug augments a simple graph by constructing a set of hyperedges $\mathcal{E}_{h}\subseteq\mathscr{P}(\mathcal{V})\setminus\varnothing$ from the original graph, via a hyperedge extraction function $h:\mathfrak{G}\times\mathbb{R}^{|\mathcal{V}|\times d_{v}}\times\mathbb{R}^{|\mathcal{E}|\times d_{e}}\to\mathscr{P}(\mathscr{P}(\mathcal{V}))$, where $\mathfrak{G}$ denotes the input graph space, $\mathbb{R}^{|\mathcal{V}|\times d_{v}}$ denotes the $d_{v}$-dimensional node feature space, ||×$\mathbb{R}^{|\mathcal{E}|\times d_{e}}$ denotes the $d_{e}$-dimensional edge feature space, and $\mathscr{P}(\mathscr{P}(\mathcal{V}))$ represents the set of all possible collections of hyperedges, which defines the output hyperedge space:
$\mathcal{E}_{h}=h(\mathcal{G},\mathbf{X},\mathbf{E})$ (1)
where $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d_{v}}$ denotes the node feature matrix of the graph, with each row $\mathbf{x}_{v}\in\mathbb{R}^{d_{v}}$ being the $d_{v}$-dimensional features of node $v$, and $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$ denotes the edge feature matrix of the graph, with each row $\mathbf{e}_{e}\in\mathbb{R}^{d_{e}}$ being the $d_{e}$-dimensional features of edge $e$. Each hyperedge $e_{h}\in\mathcal{E}_{h}$ is a subset of $\mathcal{V}$ containing at least three nodes, thereby capturing complex multi-node relations. TopoAug then constructs the combinatorial complex $(\mathcal{V},\mathcal{X},\mathrm{rk})$ from the original graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and extracted hyperedges $\mathcal{E}_{h}$, abbreviated as $\mathcal{X}$, according to the following rules:
[…]
$\mathcal{X}=\{\{v\}\>|\>v\in\mathcal{V}\}\cup\{\{u,v\}\>|\>(u,v)\in\mathcal{E}\}\cup\mathcal{E}_{h}$ (3)
$\forall x\in\mathcal{X}.\>\mathrm{rk}(x)=\begin{cases}0&\text{for}x=\{v\}\text{ where }v\in\mathcal{V}\\ 1&\text{for }x=\{u,v\}\text{ where }(u,v)\in\mathcal{E}\\ 2&\text{otherwise (i.e., for }x\in\mathcal{E}_{h}\text{)}\end{cases}$ (4)
Depending on the nature of the original graph, the virtual hyperedges can be constructed in one of the following three ways:
(a) Without TopoAug
(b) With TopoAug